Albert's father is 24 years older than Albert is. In 8 years, his father will be twice as old as Albert. How old is albert now? How old is his father?
Answer by fractalier(6550)

Call Albert and his father's ages NOW, a and f.
Then we have
f = a + 24
and
f + 8 = 2(a + 8)
Substitute for f in the second and get
a + 24 + 8 = 2(a + 8)
a + 32 = 2a + 16
and
a = 16
f = 40